Leg 173 Operations Synopsis

To Site 1066


At 1800 hr on Monday, 21 April 1997, the last line was passed ashore and the vessel headed down the river Tagus to the entrance of Lisbon harbor. At 1830 hr the harbor pilot disembarked, and the vessel got under way at full speed toward the first site of Leg 173. Clocks were set back 1 hr to GMT during the night. The 170-nmi transit to Site 1065 was accomplished at an average speed of 7.2 kt. During the transit a reduced speed was maintained because official drilling clearance had yet to be received. The vessel arrived on location at 1630 hr 22 April 97. Thrusters and hydrophones were lowered and the vessel began "drifting" on-location while in dynamic positioning (DP) mode. At 1700 hr, word was received that a beacon could be deployed, drill pipe tripped to the seafloor, and a subsea television (TV) survey conducted while awaiting clearance. Permission to spud the hole remained withheld.

SITE 1065
(Proposed Site Iberia-8B)

Hole 1065A
At 1800 hr on 22 April, a Datasonics beacon was dropped on the GPS coordinates for Site 1065 officially initiating Hole 1065A. The beacon subsequently failed within 2.25 hrs and a second beacon was deployed at 2015 hr. The precision depth recorder (PDR) indicated a seafloor depth adjusted to the rig floor of 4780.4 m. The pipe was tripped to the seafloor and the subsea camera was deployed. While waiting on clearance to spud the hole, a visual TV survey was done for 1000 m to the west and 880 m to the east of the site. The GPS coordinates placed this site ~5 nmi west of the TAGIDE 1 cable and ~10 nmi to the east of the TAGIDE 2 cable. There was no visual indication that the site location was at or near any subsea cables.

At 1715 hr 23 April 97, word was received that the hold on drilling clearance for the Portuguese EEZ had been removed by the Ministry of Foreign Affairs, and we were now free to continue with Leg 173 scientific drilling and coring operations. Within fifteen minutes, at 1730 hr, Hole 1065 was spudded at a seafloor depth of 4781.0 m. The bit was washed ahead to 40 mbsf without rotation while the subsea TV was retrieved. Once the TV was recovered, drilling operations commenced using the rotary core barrel (RCB) coring system with a center bit installed. The hole was drilled to a depth of 251.0 mbsf in ~12.5 hr, including the wireline time for one center bit inspection at a depth of 135.8 mbsf and final recovery of the center bit before initiating coring operations.

Continuous wireline coring proceeded initially with an acceptable 40.9% recovery from 251.0 to 308.8 mbsf (nannofossil chalk and claystone). Recovery became very poor (<5%) below 308.8 mbsf (soft clay, siltstone, sandstone, and conglomerate). Acoustic basement was reached at a depth of ~501.5 mbsf. At this point the rate of penetration (ROP) continued to be slow at 3.8 m/hr; however, recovery did improve slightly, averaging 13.2% over the interval from 501.5 to 631.4 mbsf. Because of the poor drilling conditions and the lack of crystalline basement beneath the reflector defining the acoustic basement, coring was halted at a depth of 631.4 mbsf at 1830 hr 29 April, and preparations were begun for wireline logging.

The hole was swept with 30 bbl of sepiolite drilling mud after Core 1065A-34R and then swept again with 30 additional bbl that was timed to reach the bit at the conclusion of the last core (1065A-35R). A wiper trip was made to a depth of 100 mbsf and back. Pipe overpull of 50k was noted at 331.0 mbsf during the trip out, and a bridge was encountered at 585.0 mbsf on the return. A minimal amount of fill (~1.0 m) was found on bottom. The bit was released, the sleeve shifted, the hole was displaced with 200 bbl of sepiolite mud, and the pipe was placed at a logging depth of 108 mbsf. Wireline logging proceeded with the first suite of logging tools, the Triple Combination, consisting of NGS (Natural Gamma-ray Sonde), DIT (Dual Induction Tool), HLDT (Hostile Litho-Density Tool), APS (Accelerator Porosity Sonde) and the Lamont temperature tool. The first run had trouble passing a tight spot in the hole at 321.0 mbsf, but the tools were eventually worked through and reached within 5 m of the bottom. On the return trip the tools would not pass the same tight spot. After 1.5 hours the tools were eventually worked through the suspected keyseat and recovered aboard ship. The pipe was then lowered across the bad zone placing the open ended pipe at a depth of 5128.0 m (347.0 mbsf). The second logging run consisted of the NGT (Natural Gamma-ray Tool), SDT (Sonic Digital Tool), and FMS. These tools reached to within 17.4 m of bottom and yielded some good logs in areas where the hole was not vastly over gauge. Upon concluding wireline logging operations, the pipe was pulled clear of the seafloor and the primary positioning beacon was released and recovered.

Back to Top


While the pipe trip continued, the vessel was moved in DP mode at ~0.5 kt to a way point located halfway between Sites 1066 and 1067. The end of the bottom hole assembly (BHA) reached the rig floor at 0830 hr 1 May 1997, officially ending Hole 1065A and beginning Hole 1066A. A new C-4 RCB bit and bit release were made up and the pipe was started back to bottom. Once the way point was reached (~375 m from each of the two sites) a primary positioning beacon was deployed at 1345 hr 1 May 97. The vessel then continued on to the GPS coordinates for Site 1066.

SITE 1066
(Proposed Site Iberia-9B)

After completing Site 1065, the vessel moved to a point halfway between Sites 1066 and 1067 where a beacon was deployed at 1345 hr on 1 May 1997. A jet-in test was done at Site 1066 to avoid an additional round trip of the drill string if a reentry cone had to be deployed at this site later. A subsea TV survey was conducted in the vicinity of the site coordinates prior to initiating the jet in test. No evidence of subsea cables was observed. Hole 1066A was spudded at 2100 hr on 1 May 1997. The seafloor was determined to be 5032.0 m (drill-pipe measurement [DPM]) and the jet-in test proceeded to a depth of 5093.0 m or 61.0 mbsf. The drill string was pulled clear of the seafloor at 2245 hr that same day ending Hole 1066A. The drill string was secured with knobbies through the guide horn and the vessel proceeded to move in DP mode the ~750 m to Site 1067.

Back to Top

To 173 Operations: Sites 1067 and 1068

To 173 Table of Contents

ODP Publications

ODP Home